# Standardizing the Measurement of Text Diversity: A Tool and a Comparative Analysis of Scores

### v1

@@ 5 Text Length as Confounder
The top panel of the table shows scores for human-written texts: the original article given as input for summarization, the baseline summary consisting of the first three sentences of the news article and the human reference summary. These scores serve as a reference point with respect to which to interpret the scores for models. The expectation is that the human texts are more diverse than those produced by LLMs, with the caveat that the texts were scraped from the web, so may contain HTML, ads, and page layout artefacts which might be repetitive (Salkar et al., 2022).
The human texts differ by length and the sources of longer texts appear to be less diverse. The association between the length of the produced texts and their diversity is similarly pronounced in the XSUM dataset, as seen Table 4. Text length as a confounder for diversity has been reported in prior work (Salkar et al., 2022), along with potential methods to adjust for this, e.g., sampling blocks of fixed size (Covington & McFall, 2010).
@@
Table 6: Correlation between each score and total number of words (concatenated text) for CNN/Daily Mail.
@@ 6 Diversity of Model Summaries
The confound of length complicates reporting. On both CNN/DM and XSUM (cf. Tables 3 and 4), StableLM produces the longest summaries. All scores indicate that these are the least diverse, most likely due to the length confound. In both sets of results, we look for models that produce shorter summaries that are less diverse. These findings are notable and hold, despite length differences.
Three types of differences are marked in the tables. Model summaries that are shorter but less diverse than human summaries are marked in bold. Human texts here are written by journalists, so the expectation is that they would be more diverse. More bold entries in a column indicate that the score captures difference between human and machine diversity, which is a desirable trait. Underlined entries mark models that are less diverse than other models that produce longer summaries. The more underlined entries there are for a model, the more indicators there are that its output is less diverse. Asterisks mark models that appear more diverse than a human text of shorter length.
The most interesting diversity scores are ones that capture differences between human and automatically produced text, without necessarily committing to an interpretation of which source is preferable. Human evaluation in future work will address this question. On the CNN/DM dataset, homogenization with BERTScore and MATTR are the two scores that detect no differences between human and model texts. BERTScore does not detect such differences on the XSUM dataset either. Compression ratio for part of speech sequences is the score that identifies the most differences between human and model-generated text. Self-repetition stands out as the only score that identifies model generated text as more diverse on the CNN/DM dataset. From this analysis, CR:POS and self-repetition emerge as prime candidates of reportable scores, while homogenization with BERTScore as perhaps not useful.
All scores detect at least one difference for a pair from the models we study. According to seven of the scores, Llama-2 generates texts that are less diverse than those from Mistral. FLAN-T5 is also marked as less diverse than StableBeluga according to four scores. Finally, four scores identify GPT-4 as less diverse than FLAN-T5; two of these are BERTScore homogenization, which we establish is perhaps not necessarily applicable and self-repetition, which marks human text as less diverse.
The XSUM dataset results in fewer notable observations. The one consistent findings is that FLAN-T5 produces that shortest and least diverse summaries, less diverse than other models and less diverse than human text. The type of input text clearly changes the behavior of the models and the diversity of text they produce.
@@ 7 Correlation Analysis
Here we present three sets of correlation analyses between (i) different diversity scores, (ii) the same diversity score across datasets, and (iii) diversity scores and standard reference-based evaluations.
Despite the large number of diversity scores in our list, they all revolve around $n$-gram repetition. It is of interest to know if they capture different or similar information. With this motivation in mind, we computed the correlations between every pair of scores, shown in Figure 1.
Compression ratio is highly to moderately correlated with other $n$-gram scores. The only weak correlations are with Self-BLEU and BERTScore homogenization. BERTScore homogenization and Self-BLEU are moderately correlated with each. Given the degenerate behavior of BERTScore homogenization on the analysis of summaries, reporting self-BLEU only is advisable. Finally, self-repetition is only moderately correlated with with other scores, so is informative to report as a standard score for diversity. The correlations are similar on the XSUM summaries (see Appendix 4), reinforcing the recommendation for the set of scores that should be used to capture diversity.
Diversity analysis on the CNN/DM and XSUM datasets did not indicate consistent system behavior. We further examine this mismatch, reporting in Table 2 the correlations between diversity score types values across the two datasets. Self-BLEU scores are almost perfectly correlated between the two datasets; they appear to not be affected by text source. The other scores are still moderately to highly correlated but as already observed, models are ranked differently. When reporting diversity, source of analyzed data also has to be taken into account, in addition to length.
Our guiding assumption is that output diversity and self-repetition are aspects of model behavior that are not captured by existing evaluation approaches. Here we directly test this assumption. We compute the system level correlation between the diversity scores and the traditional BERTScore and ROUGE evals, shown in Figure 2. The reference-based evaluations are only weakly correlated with the diversity metrics. Self-BLEU, however, is moderately anti-correlated with with both ROUGE-1 and BERTScore.
@@ 8 Truncating to Control Length
Table 7: Scores on CNN/DM summaries truncated to the length of the shortest summary for a given input.
For each input for summarization, we truncate all summaries to the length of the shortest one produced by any of the sources, as a crude method to remove the influence of length on scores. The resulting scores are directly comparable across sources, listed in Tables 4 and 5.
Compression ratio and Self-BLEU scores indicate that model-produced text is less diverse than human text. BERTScore homogenization scores barely vary across sources, further supporting the recommendation that this is not a useful score to report.
On the CNN/DM dataset, Self-BLEU indicates that Llama-2 and StableLM are the most repetitive models. Compression ratio also ranks these two models as the least diverse. The results are consistent on XSUM, but for that dataset Flan-T5 is also highly ranked and the most repetitive.
@@
9 Run-Time Considerations
When analyzing the diversity of large volumes of text, run-time considerations become relevant. Figure 3 provides insights about the feasibility of obtaining scores for large samples [^6] Run on a single NVIDIA Quadro RTX 8000 GPU.. The compression ratio scores are fast, with text compression utilities specifically optimized for speed. Self-repetition takes longer but acceptable time. Self-BLEU and BERTScore homogenization are prohibitively slow.
Figure 3: Mean run time (log-scale) on CNN/DM summaries. Run times increase with the number of text for the analysis. Even for small datasets, Self-BLEU and BERTScore homogenisation are unpractically slow.
10 Broader Applications
@@
10.1 Human Story Writing
Padmakumar et al. (2023) presented an analysis of human-written stories, where people wrote either by themselves or with the help of GPT-3 or GPT-3.5 Turbo. They find that using LLMs as writing partners leads to greater homogenization of the stories.
As reported by Padmakumar et al. (2023), we find that all diversity scores agree that people writing independently produce the more diverse texts (cf. Table 8). Here, story length is not an issue because the average length of stories in each setting are comparable: 375 words for writing without help, 372 words when writing with GPT-3 and 370 when writing with GPT-3.5.
@@ 10.1 Human Story Writing
Table 8: Diversity scores for the essays dataset. Working with the help of an LLM correlates with lower diversity.
10.2 Instruction-tuning Datasets
The quality and diversity of instructions are likely to result in more robust and capable systems (Sanh et al., 2022; Mishra et al., 2022). We analyze the diversity of five instruction-tuning datasets.
@@ 10.2 Instruction-tuning Datasets
In Table 9 we report diversity scores. Here datasets are ordered by size; we therefore expect that scores will be sorted in diminishing order in each column. Only deviations from this ordering are reportable. We provide details about the number of instructions and words in Appendix A.4.
Open Assistant instructions are remarkably diverse compared to the other datasets, and all diversity scores for it are more favorable than that for other datasets. Unnatural instructions are remarkable in the opposite direction, with outlier scores that are so much higher, they are likely not due to length entirely. We provide an analysis of the diversity scores with the length controlled in Appendix A.5.
Dataset CR (↓) CR: POS (↓)
Self-Rep. (↓)
Open Assistant 2.886 6.731 3.969
Unnatural Instructions 4.191 7.278 9.868
Alpaca 3.119 6.61 3.105
Super-NaturalInstructions 2.675 5.749 3.456
Dolly 2.578 6.214 2.935
11 Discussion and Recommendations
Our in-depth analyses reveal that compression ratio is an excellent score to report, easy to compute and strongly correlated with other scores used in past work. Compression ratio of part of speech sequences capture differences between human and model-generated text, so is also a good score to track. Self-repetition zeros in only on repetition of longer $n$-grams across generations and is only moderately correlated with compression ratios and is intuitively interpretable, ass desirable characteristics. Finally Self-BLEU is only weakly correlated with the previous three, so is a good complement score to report. In our analyses, we identified several drawbacks of BERTScore: it does not show differences between human and model-generated text and barely varies when adjusted for length. There is no good justification to report it.
Length of the analyzed text has to be reported alongside all these scores. When length differs, scores are not meaningfully comparable. Truncating and downsampling text is one way to produce a set of results that are intuitively comparable. Different random draws of the sample chosen to represent a dataset will likely differ in diversity; the selection may lead to unwarranted conclusions. Truncating texts prevents any possibility of discovering repetitive behavior towards the end of longer text. Future research into a principled solution for this problem is urgently needed.
Despite all this, we were able to glean meaningful insights about differences in diversity between human and model-produced text for summaries, essays and instructions.
12 Limitations
In the work presented here we do not explore human approaches to evaluating the diversity of collections of text. These are straightforward when the produced text is fairly short, as in judging the diversity of a set of questions generated for a given document (Sultan et al., 2020) or the diversity of possible continuation of a conversation (Tevet & Berant, 2021). Longer texts, as in the case of summaries, and larger collections, as in the case of instruction datasets are harder to judge for diversity.
An interface allowing people to explore the data, organized by stretches of repeated text ordered either by the length of the repeated string or the number of times it has been repeated, can facilitate human evaluation.
Acknowledgements
Appendix A Appendix
A.1 Examples of Repetitive Patterns
Dataset Token Repetition Text Pattern-Matched Text
GPT-3 ”In my opinion…” 41/100 PRP VBZ RB JJ TO VB 15/100
”It is also vital to discern…”
”It is very easy to construct…”
”It is largely inappropriate to try…”
”It is morally acceptable to focus…”
PRP VBP IN DT NN IN 12/100
”I don’t like the damsel in…”
”I fear that a cycle of…”
”I feel that an acknowledgement of…”
”I find that the inflection of…”
Instruct-GPT ”In my opinion…” 25/100 MD VB DT JJ NN IN 20/100
”It is important to…” 20/100 ”…can have a huge variety of…”
”Up with the news…” 15/100 ”…can have a negative effect on…”
”…can have a positive impact on…”
”…can have a sturdy framework for…”
PRP VBZ RB JJ TO VB 12/100
”It is also important to realize…”
”It is fairly common to hear…”
”It is indeed surprising to hear…”
”It is probably impossible to keep…”
”…it becomes very cringy to watch…”
Solo ”In my opinion…” 22/100 PRP VBZ JJ TO VB IN 9/100
”In my opinion, the…” 13/100 ”It is crucial to recognize that…”
”When it comes to…” 11/100 ”It is crucial to remember that…”
”In my opinion, I…” 10/100 ”It is unjustifiable to assume that…”
”It is important to acknowledge that…”
PRP VBP IN DT JJ NN 10/100
”I believe for the right person…”
”I do on a regular basis.”
”I fall into the second group.”
”I live in a small town…”
Table 10: Examples of exact text-match and repeating part-of-speech patterns in essays from Padmakumar & He (2023b). The number of times the pattern occurs is shown in parenthesis.
A.2 Summarization Prompts
Model Model Size Prompt
Llama-2 7B [TEXT] [INST] Summarize the above text. [/INST]
GPT-4 - [TEXT]. Summarize the above text.
Flan-T5 11B Summarize this article: [TEXT]
StableLM 7B [TEXT] <|<|USER|>|>Summarize the above text.
<|<|ASSISTANT|>|>
### Response:
A.3 XSUM Metrics
Figure 4: Correlation table between scores on XSUM.
A.4 Instruction Datasets, Details
@@
A.5 Instruction Datasets, Length Controlled

### v2

@@ 5 Text Length as a Confounder
One would expect that human-authored texts would be more diverse than those produced by LLMs (with the caveat that the texts were scraped from the web, and so may contain HTML and page layout artefacts which might be repetitive (Salkar et al., 2022)). The human texts differ by length and the sources of longer texts appear to be less diverse. The association between the length of the produced texts and their diversity is similarly pronounced in the XSUM dataset, as seen Table 3. Text length as a confounder for diversity has been reported in prior work (Salkar et al., 2022), along with potential methods to adjust for this, e.g., sampling blocks of fixed size (Covington and McFall, 2010).
@@
Table 5: Correlation between scores and total word counts (concatenated text) for CNN/Daily Mail.
@@ 6 Diversity of Model Summaries
The confound of length complicates reporting. On both CNN/DM and XSUM (cf. Tables 2 and 3), StableLM produces the longest summaries. All scores indicate that these are the least diverse, most likely due to the length confound. In both sets of results, we look for models that produce shorter summaries that are less diverse. These findings are notable and hold, despite length differences.
Three types of differences are marked in the tables. Model summaries that are shorter but less diverse than human summaries are marked in bold. Human texts here are written by journalists, so the expectation is that they would be more diverse. More bold entries in a column indicate that the score captures differences between human and machine diversity, a desirable trait. Underlined entries denote models that are less diverse than other models that produce longer summaries. The more underlined entries there are for a model, the more indicators there are that its output is less diverse. Asterisks mark models that appear more diverse than a human text of shorter length.
The most interesting diversity scores are those that capture differences between human and automatically produced text, without necessarily committing to an interpretation of which source is preferable. On the CNN/DM dataset, Hom. (BERT) and MATTR are the two scores that detect no differences between human and model texts. Compression ratio for part of speech sequences is the score that identifies the most differences between human and model-generated text. Self-repetition stands out as the only score that identifies model generated text as more diverse on the CNN/DM dataset. From this analysis, CR:POS and self-repetition emerge as prime candidates of reportable scores, while Hom. BERT is less useful.
@@ 7 Correlation Analysis
We present three sets of correlation analyses between (i) different diversity scores, (ii) the same diversity score across datasets, and (iii) diversity scores and standard reference-based evaluations. Despite the large number of diversity scores in our list, they all revolve around $n$-gram repetition. Do these capture different (complementary) information? To assess this we compute the correlations between all pairs of scores, reported in Figure 3.
Compression ratio is highly to moderately correlated with other $n$-gram scores. The only weak correlations are with Self-BLEU and Hom. (BERT). Given the degenerate behavior of Hom. (BERT) on the analysis of summaries, reporting Self-BLEU only is advisable. Finally, self-repetition is only moderately correlated with with other scores, so it is informative to report this as a standard diversity score. Correlations are similar on XSUM summaries (Appendix 6), reinforcing the recommendations here. Diversity analysis on the CNN/DM and XSUM datasets did not indicate consistent system behavior. We report the analysis in Table 10.
Our guiding assumption is that output diversity and self-repetition are aspects of model behavior that are not captured by existing evaluation approaches. Here we directly test this assumption. We compute the system level correlation between the diversity scores and the traditional BERTScore and ROUGE evals, shown in Figure 4. Reference-based evaluations are only weakly correlated with the diversity metrics. Self-BLEU, however, is moderately anti-correlated with with both ROUGE-1 and BERTScore.
@@ 8 Truncating to Control Length
Table 6: Scores on CNN/DM summaries truncated to the shortest summary length for a given input.
For each input for summarization, we truncate all summaries to the length of the shortest one produced by any of the sources as a crude means to remove the influence of length on scores. The resulting scores are directly comparable across sources, listed in Tables 6 and 4. Compression ratio and Self-BLEU scores indicate that model-produced text is less diverse than human text. Hom. (BERT) scores barely vary across sources, further supporting the recommendation that this is not a useful score to report. On the CNN/DM dataset, Self-BLEU indicates that Llama-2 and StableLM are the most repetitive models. Compression ratio also ranks these two models as the least diverse. The results are consistent on XSUM, but for that dataset Flan-T5 is also highly ranked and the most repetitive.
@@
9 Discussion and Recommendations
The diversity package and platform provides a useful way to analyze and visualize diversity in datasets. For reporting metrics, our in-depth analyses reveal that compression ratio is an excellent score to report, easy to compute and strongly correlated with other scores used in past work. Compression ratio of part of speech sequences capture differences between human and model-generated text. Self-repetition zeros in only on repetition of longer $n$-grams across generations, and is only moderately correlated with compression ratios. Finally Self-BLEU is only weakly correlated with the previous three, so is a good complement score to report. In our analyses, we identified several drawbacks of BERTScore: it does not show differences between human and model-generated text and barely varies when adjusted for length.
Length of the analyzed text has to be reported alongside all these scores. When length differs, scores are not meaningfully comparable. Truncating and downsampling text is one way to produce a set of results that are intuitively comparable. Different random draws of the sample chosen to represent a dataset will likely differ in diversity; the selection may lead to unwarranted conclusions. Truncating texts prevents any possibility of discovering repetitive behavior towards the end of longer text. Future research into a principled solution for this problem is urgently needed. Despite all this, we were able to glean meaningful insights about differences in diversity between human and model-produced text for summaries, essays and instructions.
Finally, diversity offers a platform and package in which researchers from a variety of domains can use to facilitate evaluation.
10 Limitations
In the work presented here we do not explore human approaches to evaluating the diversity of collections of text. These are straightforward when the produced text is fairly short, as in judging the diversity of a set of questions generated for a given document (Sultan et al., 2020) or the diversity of possible continuation of a conversation (Tevet and Berant, 2021). Longer texts, as in the case of summaries, and larger collections, as in the case of instruction datasets are harder to judge for diversity.
Acknowledgements
Appendix A Appendix
Appendix B Run-Time Considerations
When analyzing the diversity of large volumes of text, run-time considerations become relevant. Figure 5 provides insights about the feasibility of obtaining scores for large samples [^10] Run on a single NVIDIA Quadro RTX 8000 GPU.. The compression ratio scores are fast, with text compression utilities specifically optimized for speed. Self-repetition takes longer but acceptable time. Self-BLEU and Hom. (BERT) are prohibitively slow.
Figure 5: Mean run time (log-scale) on CNN/DM summaries. Run times increase with the number of text for the analysis. Even for small datasets, Self-BLEU and BERTScore homogenization are slow.
Appendix C Broader Applications
@@
Human Story Writing
Padmakumar et al. (2023) presented an analysis of human-written stories, where people wrote either by themselves or with the help of GPT-3 or GPT-3.5 Turbo. They find that using LLMs as writing partners leads to greater homogenization of the stories. As reported by Padmakumar et al. (2023), we find that all diversity scores agree that people writing independently produce the more diverse texts (cf. Table 7). Here, story length is not an issue because the average length of stories in each setting are comparable: 375 words for writing without help, 372 words when writing with GPT-3 and 370 when writing with GPT-3.5.
@@
Table 7: Diversity scores over essays. Working with an LLM correlates with lower diversity.
Instruction-tuning Datasets
The quality and diversity of instructions are likely to result in more robust and capable systems (Sanh et al., 2022; Mishra et al., 2022). We analyze the diversity of five instruction-tuning datasets: Open Assistant (Köpf et al., 2024), Super-NaturalInstructions (Wang et al., 2022), Unnatural Instructions (Honovich et al., 2023), Alpaca (Wang et al., 2023b), and Dolly (Conover et al., 2023).  [^11] We provide further details in Appendix C.6.
In Table 8 we report diversity scores. Here datasets are ordered by size; we therefore expect that scores will be sorted in diminishing order in each column. Only deviations from this ordering are reportable. We provide details about the number of instructions and words in Appendix C.6. Open Assistant instructions are remarkably diverse compared to the other datasets, and all diversity scores for it are more favorable than that for other datasets. Unnatural instructions are remarkable in the opposite direction, with outlier scores that are so much higher, they are likely not due to length entirely. We provide an analysis of the diversity scores with the length controlled in Appendix C.7.
Dataset CR (↓) CR: POS (↓)
Self-Rep. (↓)
Open Assistant 2.886 6.731 3.969
Unnatural Instructions 4.191 7.278 9.868
Alpaca 3.119 6.61 3.105
Super-NaturalInstructions 2.675 5.749 3.456
Dolly 2.578 6.214 2.935
C.1 Human Assessments of Diversity
C.2 Examples of Repetitive Patterns
Dataset Token Repetition Text Pattern-Matched Text
GPT-3 "In my opinion…" 41/100 PRP VBZ RB JJ TO VB 15/100
"It is also vital to discern…"
"It is very easy to construct…"
"It is largely inappropriate to try…"
"It is morally acceptable to focus…"
PRP VBP IN DT NN IN 12/100
"I don’t like the damsel in…"
"I fear that a cycle of…"
"I feel that an acknowledgement of…"
"I find that the inflection of…"
Instruct-GPT "In my opinion…" 25/100 MD VB DT JJ NN IN 20/100
"It is important to…" 20/100 "…can have a huge variety of…"
"Up with the news…" 15/100 "…can have a negative effect on…"
"…can have a positive impact on…"
"…can have a sturdy framework for…"
PRP VBZ RB JJ TO VB 12/100
"It is also important to realize…"
"It is fairly common to hear…"
"It is indeed surprising to hear…"
"It is probably impossible to keep…"
"…it becomes very cringy to watch…"
Solo "In my opinion…" 22/100 PRP VBZ JJ TO VB IN 9/100
"In my opinion, the…" 13/100 "It is crucial to recognize that…"
"When it comes to…" 11/100 "It is crucial to remember that…"
"In my opinion, I…" 10/100 "It is unjustifiable to assume that…"
"It is important to acknowledge that…"
PRP VBP IN DT JJ NN 10/100
"I believe for the right person…"
"I do on a regular basis."
"I fall into the second group."
"I live in a small town…"
Table 9: Examples of exact text-match and repeating part-of-speech patterns in essays from Padmakumar and He (2023b). The number of times the pattern occurs is shown in parenthesis.
C.3 Correlation Between Metrics
Self-BLEU scores are almost perfectly correlated between the two datasets; they appear to not be affected by text source. The other scores are still moderately to highly correlated but as already observed, models are ranked differently. When reporting diversity, source of analyzed data also has to be taken into account, in addition to length.
CR
CR: POS
NGD
Self-
Rep.
Hom.
(R-L)
Hom.
Self-
BLEU
MATTR HD-D
0.83 0.695 0.885 0.87 0.841 0.921 0.991 0.799 0.654
C.4 Summarization Prompts
Model Model Size Prompt
Llama-2 7B [TEXT] [INST] Summarize the above text. [/INST]
GPT-4 - [TEXT]. Summarize the above text.
Flan-T5 11B Summarize this article: [TEXT]
StableLM 7B [TEXT] <|<|USER|>|>Summarize the above text.
<|<|ASSISTANT|>|>
### Response:
C.5 XSUM Metrics
Figure 6: Correlation table between scores on XSUM.
C.6 Instruction Datasets, Details
@@
C.7 Instruction Datasets, Length Controlled